# Learning Task Representations from In-Context Learning

### v1

@@ 6 Conclusion
In this study, we examined how tasks in in-context learning (ICL) are internally represented by large language models (LLMs). Motivated by the limitations of existing representations on modalities beyond text, we critically evaluated their generalizability and validity. We explored the possibility of an automated approach to encode ICL tasks that generalizes across modalities, including functional regression tasks. To address this, we propose a principled framework, Learnable Task Vector (LTV), which applies seamlessly to both text and regression tasks, regardless of the underlying model architecture. Our approach represents ICL tasks as a weighted sum of attention head activations, optimized causally through gradient descent to enhance LLM performance. Building on recent theoretical insights into the role of positional information in reasoning tasks, we also introduce a new benchmark to evaluate task fidelity in a controlled setting.
Our findings demonstrate that LTV effectively preserves task fidelity when the model is prevented from extracting task information from the demonstrations. Ablation studies reveal that this is achieved by maintaining the distribution of the last hidden state similar to that of an optimally in-context-learned model. While we did not identify significant limitations, we acknowledge that optimizing LTV can be time- and resource-intensive due to data requirements. We believe this study provides a robust framework for analyzing task encodings and ICL capabilities, particularly for functional numeric tasks.
@@ Appendix A Extended Related Work
The ICL capabilities of LLMs were first identified in [3]. Since then, ICL has been extensively studied from various angles. The effects of different ICL prompts styles have been examined [32, 66]. ICL during inference time has also been explored through meta-learning analyses [1, 5, 55, 27, 8]. In addition, investigations into ICL task inference from a Bayesian perspective have been conducted [65, 59, 63, 68]. Lastly, the scalability of ICL across different model sizes has been examined [60, 59, 35]. While these studies primarily focus on the externally observable behaviors of models during inference and ICL, our study delves into the internal mechanisms of transformers to encode ICL tasks.
@@
Past analyses of the attention mechanism [53, 4, 54, 21, 39, 28, 15, 20] have revealed that attention weights often align with linguistic structures. However, these studies primarily focused on explaining the behavior of bidirectional architectures. Moreover, attention scores alone have not been found to fully explain the model’s outputs [18, 62, 2]. In our work, we aim to deepen the understanding of the role of multi-head self-attention in ICL. Specifically, we investigate the contribution of each attention head to the model’s internal representation of the ICL task.
@@
The components of transformers during ICL inference have been investigated to identify the origins of incorrect predictions and false statements [31, 10]. Similarly, numerous studies have adjusted attention mechanisms or activations of hidden layers during inference to steer model behavior [26, 45, 49, 40, 29]. It was observed that tokens representing labels in an ICL prompt might hold the semantic information crucial for the language model’s final prediction [57]. Moreover, it was suggested that certain neurons within pre-trained transformers are highly predictive of task labels and empirically demonstrate that these neurons encode task-specific skills [58]. A similar effort has been made to identify the critical layers where task information is stored [42]. It has been shown that critical layers vary across tasks and can even differ within the same task, such as between subtasks like English-to-French and French-to-English translation. In contrast, our work aims to develop a principled conceptualization—such as a function of the transformer’s components—that can effectively represent and automatically differentiate across a wide range of tasks, regardless of task modality. We further leverage this conceptualization to guide the language model’s behavior across diverse tasks, aligning with prior works discussed next.
@@
Efforts to represent ICL tasks typically start by deriving a task vector from the layer activations [11], referred to as the state, at the position of a dummy query’s separator token. This token separates inputs and outputs in the in-context examples using specific symbols such as $\rightarrow$”. The dummy query is strategically placed immediately before the actual query to ensure that the task vector remains independent of the query itself. This method was further refined by averaging the states at each of the $N$ separator tokens in an $N$-shot prompt, followed by optimization using inner and momentum techniques [25]. However, these approaches are limited to ICL prompts that incorporate a separator token (e.g., $\rightarrow$), restricting their broader applicability. Moreover, it has been shown that the principal direction of layer activation differences can effectively guide the ICL task [29], leading to the task representation known as In-Context Vectors (ICVs). Nonetheless, it has been recently argued that the focus should be on attention heads [48], as these are crucial for transferring information between token positions [50, 7]. Deriving from this finding, the Function Vector (FV) is computed as the sum of the outputs from a selectively chosen subset of attention heads based on an indirect metric [48], derived from causal inference literature [37]. To the best of our knowledge and based on our preliminary analyses, the most effective empirical representation of tasks in ICL are FVs. Therefore, our approach starts by deriving from [48], in contrast to methods the outlined in [11, 29]. Instead of merely using raw activations, we optimize weights assigned to these heads to enhance transformer performance in scenarios where ICL cannot be performed. Ultimately, this leads to a more formalized conceptualization that can be adapted to various models and tasks, whether functional regression or linguistic.
Appendix B Discussion: Why Shuffling Labels is not a Valid Approach for Functional Regression
We analyze the effect of label shuffling on in-context learning performance for two example tasks: linguistic antonym generation and linear regression. While label shuffling disrupts the model’s ability to infer relationships in the linguistic task, its impact on the numeric task differs fundamentally due to the task’s nature and the model’s pre-training, particularly if the model was pre-trained on similar functional regression tasks [8].
B.1 Linguistic Antonym Generation
@@
$\text{vanish}\rightarrow\text{appear},\text{ short}\rightarrow\text{tall},\text{ increase}\rightarrow\text{ ?}$
The model’s goal is to generate the antonym of “increase” based on in-context examples. Since LMs are pre-trained on vast text corpora without specific emphasis on antonym pairs, they rely heavily on these examples to infer the relationship. Shuffling the labels in the examples disrupts the pattern, preventing the model from identifying the intended relationship (antonymy) and generating the correct output. Any potential performance degradation in this case is directly linked to the blockage of ICL caused by label shuffling.
B.2 Numeric Linear Regression
@@
$[x_{1,1},x_{1,2},\dots,x_{1,m}]\rightarrow y_{1},\text{ }[x_{2,1},x_{2,2},\dots,x_{2,m}]\rightarrow y_{2}\text{ },\dots,\text{ }[x_{T,1},x_{T,2},\dots,x_{T,m}]\rightarrow\text{?},$
@@ C.1 Implementation
For all models, we use the huggingface implementations [64]. For regression tasks and GPT-2 pre-training, we rely on the code provided by the authors on GitHub [^1] https://github.com/dtsip/in-context-learning. Additionally, we use the authors’ code [^2] https://github.com/ericwtodd/function_vectors to include FV into our experiments. The remaining baseline methods—ICV, TV, and SV—were implemented by us.
@@ C.2 Experiments on Regression Tasks
We closely follow the experimental setup established by [8]. For completeness, we provide the relevant details here, with additional information available in the cited reference.
@@ C.2.1 Model
We consider the model’s prediction at the position corresponding to $x_{i}$ (i.e., the absolute position $12i-$ 1) as the prediction of $f(x_{i})$. Due to the model’s structure, this prediction relies solely on the pairs $\{x_{j},f(x_{j})\}$ for $j<i$ and $x_{i}$ itself. We disregard the model predictions at positions corresponding to $f(x_{i})$.
@@ C.2.2 Training
We train a model from scratch (i.e., no pre-trained weights are loaded) to predict $f(x_{i})$ for a given $x_{i}$, using the set of examples as reference. Each training prompt is generated by randomly sampling a function $f$ from the function class of interest, followed by sampling inputs $x_{i}$ from an isotropic Gaussian distribution $N(0,I_{m})$. The prompt is constructed as $p=\{x_{1},f(x_{1}),\ldots,x_{k},f(x_{k})\}$. For each input $i\leq k$ within a prompt, the model predictions $\hat{y}_{i}=M_{\theta}(x_{i}\mid p^{f}=\{x_{1},f(x_{1}),\ldots,x_{i-1},f(x_{i-1})\})$ are obtained, and the loss is computed across all prompt prefixes::
@@
During training, we average the loss across a batch of randomly generated prompts, each with different functions and inputs, and update the model parameters. The Adam optimizer [19] is employed and trained for a total of 500,000 steps with a batch size of 64, using a learning rate of $10^{-4}$ for all function classes and models.
Curriculum Learning
The training procedure is accelerated through curriculum learning. The model starts by observing prompt inputs $x_{i}$ within a smaller dimensional subspace and with fewer inputs per prompt. Both the subspace dimension and the number of examples are increased gradually. Specifically, all of the coordinates except the first $T_{\text{max, cur}}$ of $x_{i}$ are zeroed out by sampling prompts of size $T_{\text{cur}}$. For the function classes of linear and sparse linear functions, $T_{\text{max, cur}}=5$ and $T_{\text{cur}}=11$ are used initially, and $T_{\text{max, cur}}$ and $T_{\text{cur}}$ are increased by 1 and 2, respectively, every 2000 steps until reaching $T_{\text{max, cur}}=T_{\text{max}}$ and $T_{\text{cur}}=2m+1$. A different schedule is applied for 2-layer neural networks to accommodate the need for more inputs; starting from $T_{\text{max, cur}}=5$ and $T_{\text{cur}}=26$, $T_{\text{max, cur}}$ and $T_{\text{cur}}$ are incremented by 1 and 5 respectively, every 2000 steps until $T_{\text{max, cur}}=T_{\text{max}}$ and $T_{\text{cur}}=5m+1$.
@@ C.2.5 Optimizing the Baseline Methods
    Scaling: FV performed best with a scaling factor of 1.0, aligning with [48] for language tasks. ICV required scaling (e.g., 1.5) to counterbalance the dominance of hidden state activations and prevent degradation of the vanilla model’s performance.
@@ C.2.6 Other Baseline Methods
As explained in Appendix A, the methods proposed in [11] and [25] cannot be included in our experiments because they require in-context prompts to include a separator token (e.g., $\rightarrow$”) in each example. This requirement makes them incompatible with our functional regression dataset, as these tasks explicitly avoid tokenization and operate directly on raw floating-point numbers (see Appendix C.2 for details). Introducing separator tokens fundamentally changes the task by imposing a symbolic or tokenized representation of input-output relationships, contradicting the original setup proposed in [8]. Moreover, such methods rely on the tokenized format, bypassing the challenge of inferring functional relationships from raw numeric data, which renders the comparison invalid.
@@ C.3 Experiments on Language Tasks
We closely adhere to the experimental methods described in [48], providing all relevant details to ensure our report is self-sufficient.
@@ C.3.1 Datasets and Tasks
We provide a detailed overview of the natural language tasks used to evaluate the task formulations. References to the original data sources for each task are included, along with details on how they were refined and transformed into word pairs by [48]. The final datasets used in our evaluations are provided in the authors’ GitHub repository2.
@@
This text classification dataset consists of news headlines and the first few sentences of articles as inputs, with labels indicating the article’s category. Categories include Business, Science/Technology, Sports, and World [67].
@@
The dataset is from [34] and was further refined by [48]. Initially, all adjective, noun, and verb pairs from the dataset splits were combined, with duplicate entries removed. The dataset was then modified to include only word pairs where both words are tokenized as single tokens. This process resulted in 2,398 antonym pairs and 2,881 synonym pairs, with a vocabulary size of $|\mathcal{V}|=50,400$.
These datasets originally included multiple outputs for single inputs, e.g., “increase” $\rightarrow$ “decrease” and “increase” $\rightarrow$ “reduce.” However, handling such cases would require a more powerful model [48]. Therefore, the dataset has been simplified to ensure a one-to-one mapping between terms.
@@
This question-answering dataset requires a model to select the correct answer from five options, each labeled with a letter. The model generates the letter corresponding to the correct answer. For instance, given the question “Where is a business restaurant likely to be located?” and the options “A) town, B) hotel, C) mall, D) business sector, E) yellow pages,” the model should generate “D” [46].
@@
The Landmark-Country dataset contains entries pairing the name of a landmark with its corresponding country. The data pairs are sourced from [12].
@@
The Person-Instrument dataset contains entries pairing the name of a professional musician with the instrument they play. The data pairs are sourced from [12].
@@
The Person-Occupation dataset, sourced from [12], contains entries pairing the names of well-known individuals with their respective occupations.
@@
The Person-Sport dataset, sourced from [12], contains entries pairing the names of professional athletes with the sports they play.
@@
The Product-Company dataset, curated from [12], contains entries pairing the names of commercial products with the companies that sell them.
@@
The sentiment analysis task is based on the Stanford Sentiment Treebank (SST-2) dataset [44], which consists of movie review sentences labeled as either “positive” or “negative.” For example, an entry from this dataset might be: “An extremely unpleasant film. $\rightarrow$ negative.” We use the subset of SST-2 curated by [14], which excludes incomplete sentences and those with more than 10 words, resulting in 1,167 entries. For further details, see [14].
@@
The language translation dataset was constructed using data from [24], which pairs English words with their translations in French and Spanish. The train and test splits were merged into a single dataset and the cognates were filtered out. This process results in 4,705 pairs for English-French and 5,200 pairs for English-Spanish. The original datasets included multiple translations for some input words. These duplicates were filtered using GPT-4, following the method used for Antonym and Synonym.
@@
Our extractive tasks are based on a subset of the CoNLL-2003 English named entity recognition (NER) dataset [47], a widely used NLP benchmark for evaluating NER models. The NER task involves extracting the correct entity from a given sentence based on a specific property. For our study, we use three tasks derived by [48] from CoNLL-2003: NER-person, NER-location, and NER-organization, where the labels correspond to the name of a person, location, or organization, respectively. Each dataset is constructed by merging the CoNLL-2003 “train” and “validation” splits into a single dataset and filtering sentences to include only those with a single instance of the specified class. This reduces task ambiguity, as sentences with multiple instances of the same class could have more than one correct answer.

### v2

@@ 6 Conclusion
We investigated the internal representation of in-context learning (ICL) tasks in large language models (LLMs). We propose the Learnable Task Vector (LTV) framework, which encodes ICL tasks as a weighted sum of attention head activations optimized through causal gradient descent. LTV generalizes across both text and regression tasks, addressing the limitations of existing representations. Through empirical validation using a novel benchmark designed to analyze task encodings, we demonstrate that LTV preserves task fidelity when models fail to extract task information from demonstrations. Our findings further suggest that this preservation is achieved by maintaining the last hidden state distribution similar to that of an optimally in-context-learned model, bringing us closer to understanding how task steering works in general.
Limitations
Although not explicitly stated in the main body, our study assumes access to the models’ hidden states, limiting the scope to open- or white-box models. Additionally, we acknowledge that optimizing LTV can be time- and resource-intensive, e.g., requiring 120K iterations over 10K samples for regression. LTV training serves only to learn a fixed task representation vector while keeping model weights frozen, and it is not intended to replace ICL or few-shot prompting. A dataset for LTV training must first be generated or collected. For functional regression, this is relatively straightforward, as data and regression coefficients can be sampled from distributions (e.g., zero-mean Gaussian). However, for language data, it must be collected or generated using a generative AI tool. As a reminder, we utilized data collected by Todd et al. (2024) based on widely studied datasets such as SST-2 (Socher et al., 2013a) and CoNLL-2003 (Tjong Kim Sang and De Meulder, 2003). Finally, because hidden layer access is required, deploying an open-source model on GPUs may demand significant computational resources.
Ethical Considerations
We do not identify any direct application of our study for unethical purposes.
@@ Appendix A Extended Related Work
The ICL capabilities of LLMs were first identified by Brown et al. (2020). Since then, ICL has been extensively studied from various angles. Min et al. (2022); Yoo et al. (2022) have examined the effects of different ICL prompts styles. ICL during inference time has also been explored through meta-learning analyses (Akyürek et al., 2023; Dai et al., 2023; Von Oswald et al., 2023; Li et al., 2023b; Garg et al., 2023). In addition, investigations into ICL task inference from a Bayesian perspective have been conducted (Xie et al., 2022; Wang et al., 2023b; Wies et al., 2023; Zhang et al., 2023). Lastly, the scalability of ICL across different model sizes has been examined by Wei et al. (2023); Wang et al. (2023b); Pan et al. (2023). While these studies primarily focus on the externally observable behaviors of models during inference and ICL, our study delves into the internal mechanisms of transformers to encode ICL tasks.
@@
Past analyses of the attention mechanism (Voita et al., 2018; Clark et al., 2019; Voita et al., 2019; Kovaleva et al., 2019; Reif et al., 2019; Lin et al., 2019; Htut et al., 2019; Kobayashi et al., 2020) have revealed that attention weights often align with linguistic structures. However, these studies primarily focused on explaining the behavior of bidirectional architectures. Moreover, attention scores alone have not been found to fully explain the model’s outputs (Jain and Wallace, 2019; Wiegreffe and Pinter, 2019; Bibal et al., 2022). In our work, we aim to deepen the understanding of the role of multi-head self-attention in ICL. Specifically, we investigate the contribution of each attention head to the model’s internal representation of the ICL task.
@@
The components of transformers during ICL inference have been investigated to identify the origins of incorrect predictions and false statements (Merullo et al., 2024; Halawi et al., 2024). Similarly, numerous studies have adjusted attention mechanisms or activations of hidden layers during inference to steer model behavior (Li et al., 2023a; Subramani et al., 2022; Turner et al., 2023; Rimsky et al., 2024; Liu et al., 2024). It was observed that tokens representing labels in an ICL prompt might hold the semantic information crucial for the language model’s final prediction (Wang et al., 2023a). Moreover, it was suggested that certain neurons within pre-trained transformers are highly predictive of task labels and empirically demonstrate that these neurons encode task-specific skills (Wang et al., 2022). A similar effort has been made to identify the critical layers where task information is stored (Sia et al., 2024). It has been shown that critical layers vary across tasks and can even differ within the same task, such as between subtasks like English-to-French and French-to-English translation. In contrast, our work aims to develop a principled conceptualization—such as a function of the transformer’s components—that can effectively represent and automatically differentiate across a wide range of tasks, regardless of task modality. We further leverage this conceptualization to guide the language model’s behavior across diverse tasks, aligning with prior works discussed next.
@@
Efforts to represent ICL tasks typically start by deriving a task vector from the layer activations (Hendel et al., 2023), referred to as the state, at the position of a dummy query’s separator token. This token separates inputs and outputs in the in-context examples using specific symbols such as $\rightarrow$”. The dummy query is strategically placed immediately before the actual query to ensure that the task vector remains independent of the query itself. This method was further refined by averaging the states at each of the $N$ separator tokens in an $N$-shot prompt, followed by optimization using inner and momentum techniques (Li et al., 2024). However, these approaches are limited to ICL prompts that incorporate a separator token (e.g., $\rightarrow$”), restricting their broader applicability. Moreover, it has been shown that the principal direction of layer activation differences can effectively guide the ICL task (Liu et al., 2024), leading to the task representation known as In-Context Vectors (ICVs). Nonetheless, it has been recently argued that the focus should be on attention heads (Todd et al., 2024), as these are crucial for transferring information between token positions (Vaswani et al., 2017; Elhage et al., 2021b).
Deriving from this finding, the Function Vector (FV) is computed as the sum of the outputs from a selectively chosen subset of attention heads based on an indirect metric (Todd et al., 2024), derived from causal inference literature (Pearl, 2001). To the best of our knowledge and based on our preliminary analyses, the most effective empirical representation of tasks in ICL are FVs. Therefore, our approach starts by deriving from Todd et al. (2024), in contrast to methods the outlined by Hendel et al. (2023); Liu et al. (2024). Instead of merely using raw activations, we optimize weights assigned to these heads to enhance transformer performance in scenarios where ICL cannot be performed. Ultimately, this leads to a more formalized conceptualization that can be adapted to various models and tasks, whether functional regression or linguistic.
Appendix B Discussions of Alternative Approaches to Assess Task Representations
Section B.1 explains why shuffling labels in regression tasks does not prevent the model from extracting task information. Section B.2 discusses why probing is ineffective for assessing the precision of task representations.
B.1 Why Shuffling Labels is not a Valid Approach for Functional Regression
We analyze the effect of label shuffling on in-context learning performance for two example tasks: linguistic antonym generation and linear regression. While label shuffling disrupts the model’s ability to infer relationships in the linguistic task, its impact on the numeric task differs fundamentally due to the task’s nature and the model’s pre-training, particularly if the model was pre-trained on similar functional regression tasks (Garg et al., 2023).
B.1.1 Linguistic Antonym Generation
@@
[…]
The model’s goal is to generate the antonym of “slow” based on in-context examples. Since LMs are pre-trained on vast text corpora without specific emphasis on antonym pairs, they rely heavily on these examples to infer the relationship. Shuffling the labels in the examples disrupts the pattern, preventing the model from identifying the intended relationship (antonymy) and generating the correct output. Any potential performance degradation in this case is directly linked to the blockage of ICL caused by label shuffling.
B.1.2 Numeric Linear Regression
@@
$[x_{1,1},x_{1,2},\dots,x_{1,m}]\rightarrow y_{1},$
[…]
$\vdots$
$[x_{T,1},x_{T,2},\dots,x_{T,m}]\rightarrow\text{?},$
@@
B.2 Why Probing Would be Ineffective to Assess Task Representations
We outline a potential approach for probing experiments. One method is to assess whether the LTV encodes task information by using it to predict task-relevant quantities—in this case, quantities in regression tasks or correct labels in language tasks. This prediction is performed by training a linear predictor or a two-layer neural network.
However, we argue that even if the probe achieves high accuracy, this does not demonstrate that the LTV (or any input used for probing) truly encodes task information. Probing offers only indirect evidence; it shows that a mapping exists from the input to the target quantity, not that the input itself is a meaningful representation of the task. It simply indicates that the input contains enough information to predict the output.
In an extreme case, we could apply an arbitrary invertible transformation to the prompt and use that as input to the probe. The probe would likely achieve good performance because the prompt contains sufficient information to predict the task-relevant output. However, this input would clearly not be a meaningful representation, as it is too coarse.
In our case, the LTV is a representation specific to the LLM architecture. Moreover, we are not concerned with whether it can directly predict the task-relevant output—this is not how these vectors are intended to be used. Instead, they are applied to latent layers of the LLM to steer output generation, which is exactly how we evaluate their effectiveness.
@@ C.1 Implementation
For all models, we use the huggingface implementations (Wolf et al., 2020). For regression tasks and GPT-2 pre-training, we rely on the code provided by the authors on GitHub [^2] https://github.com/dtsip/in-context-learning. Additionally, we use the authors’ code [^3] https://github.com/ericwtodd/function_vectors to include FV into our experiments. The remaining baseline methods—ICV, TV, and SV—were implemented by us.
@@ C.2 Experiments on Regression Tasks
We closely follow the experimental setup established by Garg et al. (2023). For completeness, we provide the relevant details here, with additional information available in the cited reference.
@@ C.2.1 Model
We consider the model’s prediction at the position corresponding to $x_{i}$ (i.e., the absolute position 2i12i-1) as the prediction of $f(x_{i})$. Due to the model’s structure, this prediction relies solely on the pairs $\{x_{j},f(x_{j})\}$ for $j<i$ and $x_{i}$ itself. We disregard the model predictions at positions corresponding to $f(x_{i})$.
@@ C.2.2 Training
We train a model from scratch (i.e., no pre-trained weights are loaded) to predict $f(x_{i})$ for a given $x_{i}$, using the set of examples as reference. Each training prompt is generated by randomly sampling a function $f$ from the function class of interest, followed by sampling inputs $x_{i}$ from an isotropic Gaussian distribution $N(0,I_{m})$. The prompt is constructed as $p=\{x_{1},f(x_{1}),\ldots,x_{k},f(x_{k})\}$. For each input $i\leq k$ within a prompt, the model predictions $\hat{y}_{i}=M_{\theta}(x_{i}\mid p^{f}=\{x_{1},f(x_{1}),\ldots,x_{i-1},f(x_{i-1})\})$ are obtained, and the loss is computed across all prompt prefixes:
$\underset{\theta}{\text{min}}\text{ }\mathbb{E}_{f\sim\mathcal{D}_{\mathcal{F}},x\sim\mathcal{D}_{\mathcal{X}}}\bigg[\frac{1}{T+1}\sum_{i=0}^{T}\Big(M_{\theta}\big(p^{f,i}\big)-f(x_{i+1})\Big)^{2}\bigg],$
During training, we average the loss across a batch of randomly generated prompts, each with different functions and inputs, and update the model parameters. The Adam optimizer (Kingma and Ba, 2015) is employed and trained for a total of 500,000 steps with a batch size of 64, using a learning rate of $10^{-4}$ for all function classes and models.
Curriculum Learning
@@ C.2.3 Sampling the Functions
We also experimented with decision trees as a function class, as done by Garg et al. (2023). However, we found that decision trees exhibit less sensitivity to positional encodings. This is because their feature-based reasoning relies on discrete hierarchical rules rather than point-by-point mappings. By grouping multiple inputs under the same label, decision trees largely bypass positional dependencies and are minimally influenced by untrained positional encodings.
@@ C.2.5 Optimizing the Baseline Methods
    Scaling: FV performed best with a scaling factor of 1.0, aligning with Todd et al. (2024) for language tasks. ICV required scaling (e.g., 1.5) to counterbalance the dominance of hidden state activations and prevent degradation of the vanilla model’s performance.
@@ C.2.6 Other Baseline Methods
As explained in Appendix A, Hendel et al.’s (2023) and Li et al.’s (2024) methods cannot be included in our experiments because they require in-context prompts to include a separator token (e.g., $\rightarrow$”) in each example. This requirement makes them incompatible with our functional regression dataset, as these tasks explicitly avoid tokenization and operate directly on raw floating-point numbers (see Appendix C.2 for details). Introducing separator tokens fundamentally changes the task by imposing a symbolic or tokenized representation of input-output relationships, contradicting Garg et al.’s (2023) setup. Moreover, such methods rely on the tokenized format, bypassing the challenge of inferring functional relationships from raw numeric data, which renders the comparison invalid.
@@ C.3 Experiments on Language Tasks
We closely adhere to the experimental methods described by Todd et al. (2024), providing all relevant details to ensure our report is self-sufficient.
@@ C.3.1 Datasets and Tasks
We provide a detailed overview of the natural language tasks used to evaluate the task formulations. References to the original data sources for each task are included, along with details on how they were refined and transformed into word pairs by Todd et al. (2024). The final datasets used in our evaluations are provided in the authors’ GitHub repository3.
@@
This text classification dataset consists of news headlines and the first few sentences of articles as inputs, with labels indicating the article’s category. Categories include Business, Science/Technology, Sports, and World (Zhang et al., 2015).
@@
The dataset is Nguyen et al.’s (2017) and was further refined by Todd et al. (2024). Initially, all adjective, noun, and verb pairs from the dataset splits were combined, with duplicate entries removed. The dataset was then modified to include only word pairs where both words are tokenized as single tokens. This process resulted in 2,398 antonym pairs and 2,881 synonym pairs, with a vocabulary size of $|\mathcal{V}|=50,400$.
These datasets originally included multiple outputs for single inputs, e.g., “increase” $\rightarrow$ “decrease” and “increase” $\rightarrow$ “reduce.” However, handling such cases would require a more powerful model (Todd et al., 2024). Therefore, the dataset has been simplified to ensure a one-to-one mapping between terms.
@@
This question-answering dataset requires a model to select the correct answer from five options, each labeled with a letter. The model generates the letter corresponding to the correct answer. For instance, given the question “Where is a business restaurant likely to be located?” and the options “A) town, B) hotel, C) mall, D) business sector, E) yellow pages,” the model should generate “D” (Talmor et al., 2019).
@@
The Landmark-Country dataset contains entries pairing the name of a landmark with its corresponding country. The data pairs are sourced from Hernandez et al. (2024).
@@
The Person-Instrument dataset contains entries pairing the name of a professional musician with the instrument they play. The data pairs are sourced from Hernandez et al. (2024).
@@
The Person-Occupation dataset, sourced from Hernandez et al. (2024), contains entries pairing the names of well-known individuals with their respective occupations.
@@
The Person-Sport dataset, sourced from Hernandez et al. (2024), contains entries pairing the names of professional athletes with the sports they play.
@@
The Product-Company dataset, curated from Hernandez et al. (2024), contains entries pairing the names of commercial products with the companies that sell them.
@@
The sentiment analysis task is based on the Stanford Sentiment Treebank (SST-2) dataset (Socher et al., 2013b), which consists of movie review sentences labeled as either “positive” or “negative.” For example, an entry from this dataset might be: “An extremely unpleasant film. $\rightarrow$ negative.” We use the subset of SST-2 curated by Honovich et al. (2023), which excludes incomplete sentences and those with more than 10 words, resulting in 1,167 entries. For further details, see Honovich et al.’s (2023) study.
@@
The language translation dataset was constructed using data from Lample et al. (2018), which pairs English words with their translations in French and Spanish. The train and test splits were merged into a single dataset and the cognates were filtered out. This process results in 4,705 pairs for English-French and 5,200 pairs for English-Spanish. The original datasets included multiple translations for some input words. These duplicates were filtered using GPT-4, following the method used for Antonym and Synonym.
@@
Our extractive tasks are based on a subset of the CoNLL-2003 English named entity recognition (NER) dataset (Tjong Kim Sang and De Meulder, 2003), a widely used NLP benchmark for evaluating NER models. The NER task involves extracting the correct entity from a given sentence based on a specific property. For our study, we use three tasks derived by Todd et al. (2024) from CoNLL-2003: NER-person, NER-location, and NER-organization, where the labels correspond to the name of a person, location, or organization, respectively. Each dataset is constructed by merging the CoNLL-2003 “train” and “validation” splits into a single dataset and filtering sentences to include only those with a single instance of the specified class. This reduces task ambiguity, as sentences with multiple instances of the same class could have more than one correct answer.